Question #1: A container contains THREE RED, FIVE WHITE and EIGHT GREEN balls.
You are to draw a ball put it back and draw another ball.
What is the probability of getting TWO green balls?
Question #2: Using the same container as in problem # 9 above and selecting one ball followed by selecting another but NOT replacing the first ball what is the probability of selecting
TWO GREEN BALLS?

Answer by Mathtut(3670)

1. 8/16(8/16)=1/4 probability
2. 8/16(7/15)=7/30 probability